# The modified Physics-Informed Hybrid Parallel Kolmogorov–Arnold and Multilayer Perceptron Architecture with domain decomposition

### v1

@@ 1 Introduction
To improve the performance of PINNs in complex physical problems, researchers have proposed various improved algorithms for the PINNs. One of the key focuses is to adopt a neural network architecture with enhanced representational capabilities. Among these networks, Convolutional Neural Networks (CNNs) [14] and Graph Neural Networks (GNNs) [15] have been explored as alternatives to traditional MLP. Furthermore, an important advancement is the Kolmogorov–Arnold Network (KAN) proposed by Li et al. [17]. Based on the Kolmogorov–Arnold representation theorem, KAN expresses an arbitrary multivariate continuous function as a finite composition of univariate functions. The hierarchical function composition mechanism of KAN aligns intrinsically with the superposition principle of differential equation solutions. The mechanism not only improves the ability to approximate highly nonlinear problems but also enables seamless integration of physical priors and data-driven learning through a combination of learnable basis functions. Consequently, the adoption of KAN enhances model interpretability and expressive power for complex physical phenomena, providing a more robust and reliable framework for modeling intricate physical systems. In addition to the network architecture, another significant focus is to enhance the optimization process of PINNs, such as adaptive weighting methods [18], residual-based attention mechanisms [19], adaptive residual point selection techniques [20], causal constraint methods [21], multi-fidelity PINNs [22], progressive hierarchical learning frameworks [23], Fourier Feature Networks [25], Multi-scale Deep Neural Networks [26] and Pre-trained Physics-Informed Neural Networks (PT-PINN) [27]. These methods improve the performance and convergence by introducing additional mechanisms (such as adaptation, attention, sampling, etc.), but they increase the complexity of the algorithms and the reliance on hyperparameters or prior knowledge.
@@
As another improvement to PINNs, integrated domain decomposition technique has been successfully applied to solve problems involving multi-scale and high-frequency phenomena. PINNs with domain decomposition technique divide complex computational domains into multiple smaller subdomains for distributed solutions, which significantly enhances the computational speed, robustness, and solution accuracy of various deep learning models. Thus, the strategy of integrating various improved PINNs with domain decomposition techniques has become a computational paradigm that attracts increasing attention due to its superior convergence rate and higher computational accuracy. For example, the extended Physics-Informed Neural Network (XPINN) [28] employs a non-overlapping domain decomposition strategy, where independent neural networks are trained on separate subdomains. However, the absence of overlapping regions between subdomains introduces discontinuities at subdomain boundaries, necessitating additional coupling loss terms to facilitate information exchange between subdomains. In contrast, the Finite Basis Physics-Informed Neural Network (FBPINN) [33] adopts a more effective overlapping domain decomposition strategy, eliminating coupling loss terms and ensuring solution continuity across subdomains for better accuracy.
@@ 2.1.1 MLP and KAN
Recently, many KAN architectural variants have been proposed. In particular, the Radial Basis Function (RBF) KANs [30] and Chebyshev KANs [31] maintain compatibility with domain decomposition technique while achieving superior computational efficiency and approximation accuracy compared to conventional B-spline implementations. Based on these developments, we employ a Fourier KAN architecture that replaces B-spline basis with Fourier series, achieving enhanced performance in high-frequency component approximation compared to conventional approaches. The Fourier series are denoted as follows
@@ 2.2.1 Overlapping domain decomposition
Without loss of generality, we demonstrate the domain decomposition technique on the domain $\Omega=[0,l_{1}]\times[0,l_{2}]\times\dots\times[0,l_{d}]$, where $l_{j}(j=1,\dots,d$) represents the length in the $j$-th direction of $\Omega$. Following the framework established in [33, 35], we partition $\Omega$ into $N$ overlapping subdomains $\{\Omega_{i}\}_{i=1}^{N}$, where each $\Omega_{i}$ maintains non-empty intersection with its adjacent subdomains, that is

### v2

@@ 1 Introduction
To improve the performance of PINNs in complex physical problems, researchers have proposed various improved algorithms for the PINNs. One of the key focuses is to adopt a neural network architecture with enhanced representational capabilities. Among these networks, Convolutional Neural Networks (CNNs) [14] and Graph Neural Networks (GNNs) [15] have been explored as alternatives to traditional MLP. Furthermore, an important advancement is the Kolmogorov–Arnold Network (KAN) proposed by Li et al. [17]. Based on the Kolmogorov–Arnold representation theorem, KAN expresses an arbitrary multivariate continuous function as a finite composition of univariate functions. The hierarchical function composition mechanism of KAN aligns intrinsically with the superposition principle of differential equation solutions. The mechanism not only improves the ability to approximate highly nonlinear problems but also enables seamless integration of physical priors and data-driven learning through a combination of learnable basis functions. Consequently, the adoption of KAN enhances model interpretability and expressive power for complex physical phenomena, providing a more robust and reliable framework for modeling intricate physical systems. In addition to the network architecture, another significant focus is to enhance the optimization process of PINNs, such as adaptive weighting methods [18], residual-based attention mechanisms [19], adaptive weights method based on balanced residual decay rate [20], stacked networks [21], causal constraint methods [22], multi-fidelity PINNs [23], progressive hierarchical learning frameworks [24], Fourier Feature Networks [26], Multi-scale Deep Neural Networks [27] and Pre-trained Physics-Informed Neural Networks (PT-PINN) [28]. These methods improve the performance and convergence by introducing additional mechanisms (such as adaptation, attention, sampling, etc.), but they increase the complexity of the algorithms and the reliance on hyperparameters or prior knowledge.
@@
As another improvement to PINNs, integrated domain decomposition technique has been successfully applied to solve problems involving multi-scale and high-frequency phenomena. PINNs with domain decomposition technique divide complex computational domains into multiple smaller subdomains for distributed solutions, which significantly enhances the computational speed, robustness, and solution accuracy of various deep learning models. Thus, the strategy of integrating various improved PINNs with domain decomposition techniques has become a computational paradigm that attracts increasing attention due to its superior convergence rate and higher computational accuracy. For example, the extended Physics-Informed Neural Network (XPINN) [29] employs a non-overlapping domain decomposition strategy, where independent neural networks are trained on separate subdomains. However, the absence of overlapping regions between subdomains introduces discontinuities at subdomain boundaries, necessitating additional coupling loss terms to facilitate information exchange between subdomains. In contrast, the Finite Basis Physics-Informed Neural Networks (FBPINNs) [30] adopts a more effective overlapping domain decomposition strategy, eliminating coupling loss terms and ensuring solution continuity across subdomains for better accuracy. Building upon FBPINNs, the stacking FBPINN approach [31] integrates multifidelity stacking PINNs with time domain decomposition FBPINNs to address time-dependent multiscale problems. Furthermore, multilevel FBPINNs [32], by incorporating multilevel overlapping domain decomposition into FBPINNs, significantly enhance the accuracy and computational efficiency in solving high-frequency and multiscale problems.
@@ 2.1.1 MLP and KAN
Recently, many KAN architectural variants have been proposed. In particular, the Radial Basis Function (RBF) KANs [34] and Chebyshev KANs [35] maintain compatibility with domain decomposition technique while achieving superior computational efficiency and approximation accuracy compared to conventional B-spline implementations. Based on these developments, we employ a Fourier KAN architecture [36] that replaces B-spline basis with Fourier series, achieving enhanced performance in high-frequency component approximation compared to conventional approaches. The Fourier series are denoted as follows
@@ 2.2.1 Overlapping domain decomposition
Without loss of generality, we demonstrate the domain decomposition technique on the domain $\Omega=[0,l_{1}]\times[0,l_{2}]\times\dots\times[0,l_{d}]$, where $l_{j}(j=1,\dots,d$) represents the length in the $j$-th direction of $\Omega$. Following the framework established in [30, 32, 38], we partition $\Omega$ into $N$ overlapping subdomains $\{\Omega_{i}\}_{i=1}^{N}$, where each $\Omega_{i}$ maintains non-empty intersection with its adjacent subdomains, that is